OpenStudy (anonymous):

HELP PLZZZ!! log(4x+10)=3

OpenStudy (anonymous):

@phi

OpenStudy (kamille):

i think you made a mistake writing the problem. can you rewrite it?

OpenStudy (anonymous):

its correct

OpenStudy (kamille):

it cant be just "log" without a base

OpenStudy (anonymous):

its always 10

OpenStudy (phi):

I would assume log means log base 10

OpenStudy (anonymous):

yes it does

OpenStudy (kamille):

there must be a small number like: \[\log _{1}23\]

OpenStudy (kamille):

lg is log_10(10)

OpenStudy (anonymous):

what phi said its log base 10

OpenStudy (phi):

you can "undo" the log log(4x+10)=3 can be written as 4x + 10 = 10^3 10^3 is 1000 so the problem is 4x + 10= 1000 can you finish ?

OpenStudy (kamille):

@Phi don't you use "lg" instead of log_10(10)?

OpenStudy (anonymous):

247.5?

OpenStudy (phi):

lg seems to be a newer "standard" way of writing log base 10. But log (without a base) most often means log base 10. To add to the confusion, sometimes you see lg referring to log base 2.

OpenStudy (kamille):

oh, it is so different all over the world. Do you know what is ln?

OpenStudy (phi):

to solve 4x + 10= 1000 add -10 to both sides 4x + 10 -10 = 1000 -10 simplify: on the left 10-10 is 0, and on the right 1000-10 is 990. you get 4x = 990 divide both sides by 4 x = 990/4 = 247.5 you can check this number in the original equation log(4*247.5+10)=3 and see if you really do get 3

OpenStudy (phi):

@Kamille See http://en.wikipedia.org/wiki/Logarithm#Particular_bases in particular, the column labeled "other notations" for common variations on how to abbreviate log to some base.